what is the lateral and surface area of a regular pyramid with legs the equal 5 and a one base side that equals 6
Answer by venugopalramana(3286)

what is the lateral and surface area of a regular pyramid with legs the equal 5 and a one base side that equals 6
DO YOU MEAN LATERAL EDGES AS 4 LEGS OF 5 ...AND BASE IS SQUARE OF SIDE 6...IF SO
SQUARE BASE AREA=6*6=36
EACH LATERAL SURFACE IS A TRIANGLE OF 5,5,6 SIDES
PERIMETER=2S=5+5+6=16
S=8
AREA=SQRT(S(S-A)(S-B)(S-C))=SQRT{8(8-5)(8-5)(8-6)}=SQRT(8*3*3*2)=12
4 LATERA TRIANGLES AREA =4*12=48
TSA=48+16=64